OpenStudy (anonymous):

Write a recursive formula for the sequence 7, 13, 19, 25, 31 then find the next term in the sequence?

OpenStudy (anonymous):

13-7=6 19-13=6 hence x+6 hence 31+6

OpenStudy (kamille):

well, you know that \[a _{n}=a _{1}+(n-1)d\] you can find \[d=a _{2}-a _{1}\] you know, that \[a _{1}=7 \] \[a _{2}=13\] what is d?

OpenStudy (anonymous):

is it: an = an– 1 + 6, where a1 = 7; 37?

OpenStudy (anonymous):

I just wanted to make sure the answer I got was right. ^^;;

OpenStudy (kamille):

the answer should be \[a _{_{n}}=7+6n-6=1+6n\] when n is a number of member in sequance

OpenStudy (anonymous):

no 7+(6-1)n 7+5n in this case n=6 7+30 37

OpenStudy (kamille):

for example, you need to find \[a _{6}=1+6*6=37\]

OpenStudy (anonymous):

My answers have to look like this, that's what confused me. an = an– 1 + 6, where a1 = 7; 37 an = an– 1 + 6, where a1 = 37; 7 an = an– 1 – 6, where a1 = 6; –23 an = an– 1 – 6, where a1 = 7; –8

OpenStudy (anonymous):

It's multiple choice. =/

OpenStudy (anonymous):

The first answer is correct

OpenStudy (anonymous):

That's what I figured, thanks

OpenStudy (anonymous):

for multiple choice questions approach it in the simplest way possible which is the first one both the methods are correct